# Unsourced Random Access: A Recent Paradigm for Massive Connectivity

### v1

@@ VI-B Robust URA
Looking at the URA literature, we can observe that almost all existing URA schemes are sensitive to the choice of parameters in their proposed algorithm. That is, the proposed algorithms need to be optimized every time a system hyper-parameters, such as active user load or the target PUPE changes, in order to have an acceptable performance. This makes their standardization and practical employment challenging. Therefore, modifications are needed to make the URA solutions robust against changes in system hyper-parameters.
Another robustness aspect is due to the imperfections in the assumed models or employed communication blocks. For instance, the presence of hardware impairments where the employed hardware is not as ideal as assumed, e.g., due to non-linearities of the power amplifiers or phase errors in phase-shifters, which is very probable in many URA applications where the devices are actually cheap and low-precision sensors. Only one work in the URA literature has addressed this issue [136], where a robust solution is presented for the case of residual hardware impairments in which the remaining hardware impairments after compensations are statistically modeled as additive and multiplicative noises. Therefore, the analysis of different hardware impairments on the performance of the existing URA solutions and providing robust solutions for them is an interesting research direction.
@@ VI-C Machine Learning-Aided URA
Current solutions look at the URA problem in a single shot. However, the transmission is continuous and spans several frames in many practical applications. In these cases, machine learning techniques such as reinforcement learning can be used to adapt to the environment and learn suitable transmission strategies. Machine learning techniques can also be used to simplify some blocks in the decoding process, such as user activity detection (similar to the deep learning based solutions in [148]) and channel estimation. For instance, in [149], the authors show that employing a dictionary learning based solution obviates the need for pilots for channel estimation in the fading URA scenarios while maintaining an acceptable performance. Machine learning techniques can also be utilized to make the system more robust against changes in the environment or system hyper-parameters such as the active user load, especially as the design parameters of many existing URA solutions heavily depend on these hyper-parameters.
It can be vice versa as well: URA techniques can be employed to better realize some machine learning based solutions. For instance, in federated learning, the receiver is only interested in the sum of the transmitted signals (a function of neural network weights or their updates), not them separately. The users transmit simultaneously in over-the-air federated learning, which exploits the additive nature of the wireless channel. This alongside the huge number of potential users in envisioned federated learning applications, makes URA a suitable framework for the physical layer [150, 151]. Note that in this case, PUPE is not a suitable metric as the receiver is not interested in the transmitted messages themselves but only in their summation. So, new metrics (even higher-level machine learning based ones such as accuracy) need to be employed to better capture the system performance.
@@ VI-D Integrated Sensing and Communication in URA
In [153], the problem of integrating sensing and communication is considered in a URA setup, where a massive number of unsourced and sporadically active users are connected to the system. In the resulting system, called unsourced ISAC (UNISAC), users are divided into two main categories: Sensing and communication users. The system aims to decode signals from communication users while detecting signals from active sensing users and estimating their AOAs. The authors derive an achievable bound for the UNISAC to show its efficacy in sensing the sensing users as well as in communicating with the communication ones in a URA setup. However, there are still many aspects of an ISAC system in a URA setup that need to be investigated. For instance, the problem becomes more challenging when the sensing users are passive in contrast to the active assumption in [153].
@@ VI-E URA with Feedback
In the URA literature, it is assumed that there is no coordination between the users and the receiver. However, there can be some level of coordination as long as the signal sent from the receiver is not user-specific as there is no user identity in URA. For instance, the receiver can broadcast a common pilot/beacon to the users so that they can estimate their channel coefficients or at least their large-scale components depending on their dedicated time slots. Just note that the time dedicated to this operation needs to be considered as extra overhead when comparing with the solutions without feedback. This issue is partly tackled in [154, 155, 156]. In [154], hashes of the messages of successfully decoded users and their estimated channel coefficients are fed back using beamforming techniques; while in [155], with the help of a designed feedback beacon, the users estimate their channels and infer whether they were successfully decoded using a pre-defined threshold. One application of URA can be in sensor networks where the devices partially observe a common state, which is the main interest of the receiver. For such an application, the authors in [156] devise a mechanism where the receiver broadcasts a reliability indicator of the estimated state after each transmission, using which each device can decide whether to transmit in the next iteration.
@@ VI-F RIS-Aided URA
In URA, due to the massive number of potential users, it is very probable that the line-of-sight link between some users and the receiver is either not present or heavily attenuated. In this case, employing relaying techniques such as RIS can play an important role in improving the overall system performance by providing a substitute communication link with potentially many degrees of freedom. Only a few works have considered the RIS-aided URA scenario [134, 63], which both assume a conventional passive RIS.
@@
For moving forward, a few important points can be mentioned. Firstly, there are many promising RIS configurations and types that need to be investigated in a URA setup, such as active or hybrid RIS [158] and metasurface-based RIS [159], in order to find suitable candidates. Secondly, the employment of massive arrays such as extra large massive MIMO and RIS leads to some active users being in the near field of these arrays, which makes hybrid near field-far field analysis a necessity. This is in contrast with the existing URA literature which are all in the far field regime. Thirdly, channel estimation overhead is already one of the main issues in RIS systems due to the high number of channel coefficients per user as a result of the high number of RIS elements and potentially massive number of antennas at the receiver (as the benefits of a RIS system is amplified when combined with a massive MIMO receiver). This problem is more pronounced in a RIS-aided URA system as there is a potentially massive number of even active users, which renders most existing channel estimation schemes in RIS infeasible.
VI-G Multi-Class URA
The common assumption in the URA literature is that all the users have the same number of bits to transmit or require the same quality of service. However, this may be impractical in situations where the devices have different payload requirements or power budgets, especially considering their massive number. This problem, which is called multi-class URA is addressed in only a few works in the literature [160, 161] in the coded compressed sensing framework. Hence, a further investigation of multi-class URA in different frameworks can be interesting.
VI-H Standardization
@@
There has not been any effort in the standardization of URA, neither for incorporating it in the next updates of the current protocol nor for future generations such as 6G. This requires a thorough investigation and analysis of the current competing solutions, and finding a good candidate that preferably can support both the current protocols, perhaps with small modifications, and the high demands of the future ones. For instance, in [162], with a small modification in the 5G new radio (5G-NR) release 16 protocol, the authors achieve an improvement in energy efficiency of about 5 to 10 dB compared to the existing 5G-NR solution for systems with even a moderate number of active users (10 to 50 users), employing an IDMA-based URA solution.
@@
There are two main challenges in incorporating URA in the current protocols such as 5G-NR. The first one is the small number of available preamble sequences in 5G-NR as opposed to the massive number of preamble sequences (pilots) needed in the preamble/pilot-based URA solutions. The second one is the limited number of access patterns that can be realized which is vital in transmission pattern-based URA solutions such as ODMA-based ones. This shows that one important step in the realization of URA in future protocols is to considerably increase the available pilot or transmission patterns.

### v2

@@ VI-B Robust URA
Like any other communication system, the performance of a URA system is affected by the presence of imperfections in the assumed models or employed communication blocks. For instance, the presence of hardware impairments where the employed hardware is not as ideal as assumed, e.g., due to non-linearities of the power amplifiers or phase errors in phase-shifters, which is very probable in many URA applications where the devices are actually cheap and low-precision sensors. Only one work in the URA literature has addressed this issue [126], where a robust solution is presented for the case of residual hardware impairments in which the remaining hardware impairments after compensation are statistically modeled as additive and multiplicative noises. Therefore, the analysis of different hardware impairments on the performance of existing URA solutions and providing robust solutions for them is an interesting research direction.
URA-specific challenges: Looking at the URA literature, we can observe that the URA schemes are more sensitive than the conventional ones to the choice of design parameters. That is, in order for the URA schemes to have acceptable performance, their design parameters need to be optimized every time a system hyper-parameter, such as active user load or the target PUPE, changes. This makes their standardization and practical employment challenging. Therefore, modifications are needed to make the URA solutions robust against changes in system hyper-parameters. One way is to employ stochastic optimization techniques, e.g., optimizing the expected value of the cost metric (e.g., PUPE) over the statistics of the intended system hyper-parameters in order to obtain the set of design parameters that work well in a wide range of system hyper-parameters.
@@ VI-C Machine Learning-Aided URA
Current solutions look at the URA problem in a single shot. However, the transmission is continuous and spans several frames in many practical applications. In these cases, machine learning (ML) techniques such as reinforcement learning can be used to adapt to the environment and learn suitable transmission strategies. ML techniques can also be used to simplify some blocks in the decoding process, such as user activity detection (similar to the deep learning based solutions in [148]) and channel estimation. For instance, in [149], the authors show that employing a dictionary learning based solution obviates the need for pilots for channel estimation in the fading URA scenarios while maintaining an acceptable performance. URA-specific challenges: Apart from general challenges in implementing ML solutions in communication systems, such as energy and computational constraints and compatibility with legacy systems, there are some that are specific or more pronounced in a URA setup, which is mostly due to the massive number of potential users (up to the order of millions) and their sporadic activity. Training a neural network with an input of very large size (due to the massive number of potential users) with non-stationary characteristics (due to the sporadic activity) is a very challenging task.
It can be vice versa as well: URA techniques can be employed to better realize some ML-based solutions. For instance, in federated learning (FL), the receiver is only interested in the sum of the transmitted signals (a function of neural network weights or their updates), not them separately. The users transmit simultaneously in over-the-air FL [150], which exploits the additive nature of the wireless channel. This, alongside the huge number of potential users in envisioned FL applications, makes URA a suitable framework for the physical layer [151, 152]. Challenges of URA in FL: Note that in distributed function computation applications, such as FL, where the receiver is not interested in the transmitted messages themselves but only in a function of them (e.g., their summation in FL), PUPE is not a suitable metric for the URA implementation, as it is defined on the individual messages. So, new metrics (even higher-level ML-based ones such as accuracy) need to be employed to better capture the system performance.
@@ VI-D Integrated Sensing and Communication in URA
In [154], the problem of integrating sensing and communication is considered in a URA setup, where a massive number of unsourced and sporadically active users are connected to the system. In the resulting system, called unsourced ISAC (UNISAC), users are divided into two main categories: Sensing and communication users. The system aims to decode signals from communication users while detecting signals from active sensing users and estimating their AOAs. The authors derive an achievable bound for the UNISAC to show its efficacy in sensing the sensing users as well as in communicating with the communication ones in a URA setup.
There are still many aspects of an ISAC system in a URA setup that need further investigation. For instance, the problem becomes more challenging when sensing users are passive, as opposed to the active assumption in [154]. In this case, only communication users transmit signals to the BS, with the same signal opportunistically used to sense passive targets.
URA-specific challenges: In a URA system, a massive number of users transmit sporadically without being registered or authenticated by the network. Therefore, there is no control over transmitter characteristics such as location, signal structure, and beam direction, which are critical for reliable passive target sensing. Additionally, the random access nature of user transmissions causes heavy interference, a phenomenon uncommon in traditional sensing techniques, resulting in a lack of solutions to mitigate it. Therefore, advanced algorithms are needed to address these challenges and enhance the system’s ability to sense passive targets effectively.
@@ VI-E URA with Feedback
In communication systems, feedback has been shown to greatly simplify transmission and reception mechanisms and, in some cases, even improve system performance [155]. The main challenge in employing feedback-based techniques in URA is the lack of coordination between the users and the receiver due to the massive number of potential users. However, even in URA, there can be some level of coordination as long as the signal sent from the receiver is not user-specific, as there is no user identity in URA. For instance, the receiver can broadcast a common pilot/beacon to the users so that they check if they are in the set of successfully decoded users in the previous round, and estimate their channel coefficients or at least their large-scale components depending on their dedicated time slots. Just note that the time dedicated to this operation needs to be considered as extra overhead when comparing with the solutions without feedback. This issue is partly tackled in [156, 157, 158]. In [156], hashes of the messages of successfully decoded users and their estimated channel coefficients are fed back using beamforming techniques; while in [157], with the help of a designed feedback beacon, the users estimate their channels and infer whether they were successfully decoded using a pre-defined threshold. One application of URA can be in sensor networks where the devices partially observe a common state, which is the main interest of the receiver. For such an application, the authors in [158] devise a mechanism where the receiver broadcasts a reliability indicator of the estimated state after each transmission, using which each device can decide whether to transmit in the next iteration.
@@ VI-F RIS-Aided URA
In URA, due to the massive number of potential users, it is very probable that the line-of-sight link between some users and the receiver is either not present or heavily attenuated. In this case, employing relaying techniques such as RIS can play an important role in improving the overall system performance by providing a substitute communication link with potentially many degrees of freedom. However, the huge channel estimation overhead is the main challenge in the use of RIS in URA systems. The channel estimation overhead is already one of the main issues in RIS systems due to the high number of channel coefficients per user as a result of the high number of RIS elements and the potentially massive number of antennas at the receiver (as the benefits of a RIS system are amplified when combined with a massive MIMO receiver). This problem is more pronounced in a RIS-aided URA system as there is a potentially massive number of even users, which renders most existing channel estimation schemes in RIS infeasible. There are a few works that have considered the RIS-aided URA scenario [123, 124, 125]. Although these works make the channel estimation feasible in a RIS-aided URA setup by employing an SIC-based iterative approach and time frame slotting strategies, the high computational complexity still remains a challenge.
@@
Moreover, there are many promising RIS configurations and types that need to be investigated in a URA setup, such as active or hybrid RIS [160] and metasurface-based RIS [161], in order to find suitable candidates. Moreover, the employment of massive arrays such as extra large massive MIMO and RIS leads to some active users being in the near field of these arrays, which makes hybrid near field-far field analysis a necessity. This is in contrast with the existing URA solutions, which are all in the far-field regime.
VI-G Standardization
@@
There has not been any effort in the standardization of URA, neither for incorporating it in the next updates of the current protocol nor for future generations such as 6G. This requires a thorough investigation and analysis of the current competing solutions, and finding a good candidate that preferably can support both the current protocols, perhaps with small modifications, and the high demands of the future ones. For instance, in [162], with a small modification in the 5G new radio (5G-NR) release 16 protocol, the authors achieve an improvement in energy efficiency of about 5 to 10 dB compared to the existing 5G-NR solution for systems with a small number of active users (10 to 50 users), employing an IDMA-based URA solution. However, this solution is not scalable to higher active user loads.
@@
URA-specific challenges: There are two main challenges in incorporating URA in current protocols, such as 5G-NR and beyond, for reasonable active user loads (e.g., in the order of hundreds). The first is the small number of preamble sequences available in 5G-NR as opposed to the massive number of preamble sequences (pilots) needed in the preamble/pilot-based URA solutions. The second is the limited number of access patterns that can be realized, which is vital in transmission pattern-based URA solutions such as ODMA-based ones. This shows that an important step in the realization of URA in future protocols is to considerably increase the available pilot or transmission patterns.